### v1

# A heralded single-photon source based on superpositions of squeezed states

Abstract
We propose a heralded single-photon source based on injecting a superposition of oppositely squeezed states onto a beam splitter. Our superposition of squeezed states is composed of only even photon number states (the number of photons is equal to $2,6,10,...$) meaning the probability for an emitted single photon given a heralded singe photon events is higher than what one can achieve from the usual two-mode squeezed state. This enables one to realize an enhanced heralded single-photon source. We discuss how to create this superposition of squeezed states utilizing single-mode squeezed states and the cross-Kerr nonlinearity. Our proposed method significantly improves the probability of emitting the heralded single photon compared to spontaneous parametric down conversion.
I Introduction
Single photon sources are essential tools in many quantum information-based technologies including quantum key distribution protocols [1] and linear optical quantum computation [2, 3]. As such there has been significant world-wide effort to propose and demonstrate on-demand single photon emitters. Quantum dots [4, 5, 6, 7, 8], nitrogen-vacancy centers in diamond [9, 10, 11], and negative silicon vacancy centers in diamond [12, 13, 14] have been proposed as solid state devices for realizing such emitter. Further cavity QED systems have been considered as a triggered single-photon source [15, 16, 17, 18] but the current work horse is still spontaneous parametric down conversion (SPDC) where we trigger off the idler photon [19, 20, 21]. However, the photon pair emission rate is low [$4\times 10^{-6}$ per pump photon using a periodically poled lithium niobate (PPLN) waveguide [22]].
The essence of the heralded single-photon source is the generation of pairs of photons and the reason why we use SPDC. A more efficient method for generating entangled photons is the injection of squeezed light into a 50-50 beam splitter [23]. The resulting two-mode squeezed state source can in principle generate pairs of photons with a much higher success probability. There is however still room for improvement.
In this paper, we introduce a new and more efficient approach based on superposition of squeezed state. We investigate how to generate a pair of photons by injecting an odd superposition of squeezed states onto a 50-50 beam splitter [24]. In particular we show that we can create this superposition using a single-mode squeezed state source and a cross-Kerr nonlinear medium although its heralded success probability is less than $1/2$ [25]. Finally, we determine the second-order intensity correlation functions $g^{(2)}(0)$ for the single-photon emissions of the superposition of oppositely squeezed states and the two-mode squeezed state. We use the measure $g^{(2)}(0)$ to evaluate whether or not the physical properties of a given photon source are close to those of an on-demand single-photon source [26].
This article is organized as follows. In Sec. II, we begin by discussing the generation of a pair of photons using our superposition of oppositely squeezed states determining in Secs. II.1 and II.2 the probability and efficiency of emission of the heralded single photon source. Then Sec. II.3, we consider errors induced by imperfect operations of devices. Next in Sec. III, we compare our proposed heralded source with that from the pure two-mode squeezed state. Now in Sec. IV, we determine the second-order intensity correlation functions for our source using imperfect single-photon detectors. Finally in Sec. V we provide a brief concluding discussion.
II Photon pair generation from a superposition of oppositely squeezed states
Let us begin by discussing our scheme for producing the superposition of the oppositely squeezed states using a nonlinear cross-Kerr medium. This is depicted in Fig. 1. First, we prepare a two-mode initial state where the first and second modes are given by a squeezed state $|r\rangle=\hat{S}(r)|0\rangle$ and a coherent state $|\alpha\rangle$, respectively, as
$|r\rangle_{1}=\frac{1}{\sqrt{\cosh r}}\sum_{n=0}^{\infty}\frac{\sqrt{(2n)!}}{2^{n}n!}(-\tanh r)^{n}|2n\rangle_{1},$ (1)
$|\alpha\rangle_{2}=\exp(-|\alpha|^{2}/2)\sum_{n=0}^{\infty}\frac{\alpha^{n}}{\sqrt{n!}}|n\rangle_{2},$ (2)
where $\hat{S}(r)$ is the squeezing operator and $\{|n\rangle:n=0,1,2,...\}$ represent the photon number states. Both of these states then interact with a nonlinear cross-Kerr interaction of the form $\hat{H}_{12}=\hbar\kappa\hat{a}_{1}^{\dagger}\hat{a}_{1}\hat{a}_{2}^{\dagger}\hat{a}_{2}$, where $\hat{a}_{1}$ and $\hat{a}_{2}$ are annihilation operators of modes $1$ and $2$, respectively. Formally the squeezing operator can be written as $\hat{S}(r)=\exp[(-r/2)(\hat{a}^{\dagger 2}-\hat{a}^{2})]$. Our initial state $|\Psi_{\mbox{\scriptsize in}}\rangle_{12}=|r\rangle_{1}|\alpha\rangle_{2}$ evolves over a time $\tau$ to
$|\Psi_{\mbox{\scriptsize out}}\rangle_{12}$ $=$ $\frac{1}{\sqrt{\cosh r}}\sum_{n=0}^{\infty}\frac{\sqrt{(2n)!}}{2^{n}n!}(-\tanh r)^{n}$ (3)
×$\times|2n\rangle_{1}|\alpha e^{-i2n\kappa\tau}\rangle_{2}.$
Setting $2\kappa\tau=\pi$ we have
$|\Psi_{\mbox{\scriptsize out}}\rangle_{12}$ $=$ $\frac{1}{2}{\cal N}_{+}(r)^{1/2}|r;+\rangle_{1}|\alpha\rangle_{2}$ (4)
$+\frac{1}{2}{\cal N}_{-}(r)^{1/2}|r;-\rangle_{1}|-\alpha\rangle_{2},$
where
$|r;\pm\rangle=\frac{|r\rangle\pm|-r\rangle}{\sqrt{{\cal N}_{\pm}(r)}},$ (5)
with
${\cal N}_{\pm}(r)=2\left(1\pm\frac{1}{\cosh r\sqrt{1+\tanh^{2}r}}\right).$ (6)
The states $|r;+\rangle$ and $|r;-\rangle$ are superpositions of the photon number states $\{|2n\rangle:n=0,2,4,...\}$ and $\{|2n\rangle:n=1,3,5,...\}$, respectively.
Next, distinguishing $|\alpha\rangle_{2}$ and $|-\alpha\rangle_{2}$ by homodyne measurement, we project our system to either $|r;+\rangle$ or $|r;-\rangle$ with probability ${\cal N}_{\pm}(r)/4$, respectively. In general, $|\alpha\rangle$ and $|-\alpha\rangle$ are never perfectly orthogonal but $|\langle\alpha|-\alpha\rangle|^{2}\leq 0.0012$ for $|\alpha|>1.3$. Here, we assume that $|\alpha|$ is sufficiently large. With the squeezing parameter $r=0.725$ (corresponding to a squeezing level of $6.3$ dB), ${\cal N}_{+}(r)/4$ and ${\cal N}_{-}(r)/4$ are equal to $0.833$ and $0.167$, respectively. It is worth mentioning that this probability ${\cal N}_{-}(r)/4$ monotonically increases from 00 to $1/2$ as $r$ goes to infinity.
In our scheme, we assume that we have a stable source of squeezed states $|r\rangle$, as the generation of squeezed states is a mature technology. A typical squeezing level of squeezed light used for experiments is given by $6.3$ dB which corresponds to the squeezing parameter $|r|=0.725$ [27]. In the current paper, we typically use this parameter for showing examples but note that the detection of a $15$ dB squeezed state of light was demonstrated experimentally [28].
Figure 1: Schematic illustration of a heralded single photon source using squeezed state and a cross-Kerr medium.
II.1 Heralded single-photon probability and efficiency
Now let us establish the probability of successfully emitting our heralded single photon. We input the superposition of oppositely squeezed states $|r;\pm\rangle_{a}$ into one port of a 50-50 beam splitter with the vacuum on the second as depicted in Fig. 1. The incident squeezed field is transformed to
$|r;\pm\rangle_{a}|0\rangle_{b}$$\to$ ${\cal N}_{\pm}(r)^{-1/2}[\hat{S}_{ab}(-\frac{r}{2})\hat{S}_{a}(\frac{r}{2})\hat{S}_{b}(\frac{r}{2})$ (7)
±$\pm\hat{S}_{ab}(\frac{r}{2})\hat{S}_{a}(-\frac{r}{2})\hat{S}_{b}(-\frac{r}{2})]|0\rangle_{a}|0\rangle_{b},$
where $\hat{S}_{ab}(\pm r/2)=\exp[(\pm r/2)(\hat{a}^{\dagger}\hat{b}^{\dagger}-\hat{a}\hat{b})]$ are the origins of the entanglement between the modes.
Next the probability we can detect $n_{a}$ photons at one output port and $n_{b}$ at the other is given by
$P(n_{a},n_{b};r;\pm)$ (8)
$=$ ${\cal N}_{\pm}(r)^{-1}|\langle n_{a},n_{b}|\hat{S}_{ab}(-\frac{r}{2})\hat{S}_{a}(\frac{r}{2})\hat{S}_{b}(\frac{r}{2})|0\rangle_{a}|0\rangle_{b}$
±$\pm\langle n_{a},n_{b}|\hat{S}_{ab}(\frac{r}{2})\hat{S}_{a}(-\frac{r}{2})\hat{S}_{b}(-\frac{r}{2})|0\rangle_{a}|0\rangle_{b}|^{2}.$
In contrast, if we input an initial state $|r\rangle_{a}|0\rangle_{b}=\hat{S}_{a}(r)|0\rangle_{a}|0\rangle_{b}$ onto the beam splitter, the probability is
$P(n_{a},n_{b};r)=|\langle n_{a},n_{b}|\hat{S}_{ab}(-\frac{r}{2})\hat{S}_{a}(\frac{r}{2})\hat{S}_{b}(\frac{r}{2})|0_{a},0_{b}\rangle|^{2}.$ (9)
We can then define our generation efficiencies $E(r;\pm)$ and $E(r)$ as
$E(r;\pm)=P(1,1;r;\pm)/\sum_{n=0}^{\infty}P(1,n;r;\pm).$ (10)
$E(r)=P(1,1;r)/\sum_{n=0}^{\infty}P(1,n;r).$ (11)
If we observe |±$|\pm\alpha\rangle_{2}$ with the homodyne measurement, inject $|r;\pm\rangle_{a}$ into the beam splitter, and detect the heralding single photon in one port, we then obtain a single photon from the other port with probability $E(r;\pm)$. Similarly if we inject $|r\rangle_{a}$ into the beam splitter and detect the heralding single photon, we obtain a single photon with probability $E(r)$.
II.2 Ideal single photon detection
The photon after the beam splitter with $(n_{a},n_{b})=(1,1)$ can be considered as a heralded single photon. When we generate $|r;\pm\rangle_{1}$ from $|r\rangle_{1}$, the probability that we obtain $|r;\pm\rangle_{1}$ is given by ${\cal N}_{\pm}(r)/4$. Thus, the actual probability for obtaining $n_{a}$ and $n_{b}$ photons from $|r;\pm\rangle_{1}$ is given by ${\cal N}_{\pm}(r)P(n_{a},n_{b};r;\pm)/4$.
In Fig. 2, we plot the probability $P(1,n;r)$ and $P(1,n;r;\pm)$ that we generate an $n$ photon state $|n\rangle$ in one mode and a heralding photon in the other for the input states $|r\rangle_{a}$ and $|r;\pm\rangle_{a}$ with $r=0.725$. Looking at Fig. 2, we note that $P(1,1;r;-)=0.453$, $P(1,5;r;-)=7.85\times 10^{-3}$, and $P(1,n;r;-)=0$ for $n=0,2,3,4$. Moreover, we obtain $E(r;-)=0.983$. If we generate the superposition of the oppositely squeezed states $|r;-\rangle_{a}$, it should allow us to realize the heralded single-photon source with a high probability and so be a good alternative to the SPDC. By contrast, if we prepare the squeezed state $|r\rangle_{a}$, we obtain $P(1,1;r)=7.54\times 10^{-2}$ with $E(r)=0.859$. Looking at the plot of $|r;-\rangle$, the following question comes to mind. There must be an optimal value of $r$ to maximize the probability. Answers to this question are shown in Fig. 3.
Figure 2: Plot of the probability $P(1,n;r)$ and $P(1,n;r;\pm)$ we generate an $n$ photon state $|n\rangle$ in one mode and a heralding photon in the other for the injected states $|r\rangle_{a}$ and $|r;\pm\rangle_{a}$ with $r=0.725$. The dotted blue, solid red, and dashed purple line graphs represent $P(1,n;r)$, $P(1,n;r;-)$, and $P(1,n;r;+)$, respectively. The line graph of $P(1,n;r;+)$ overlaps the horizontal axis of $n$ mostly.
Figure 3: (a) Plot of the probability ${\cal N}_{-}(r)P(1,1;r;-)/4$ we obtain a heralded single photon by injecting $|r;-\rangle_{a}$ for $0\leq r\leq 2$. When $r=1.146$, the graph reaches its maximum value 0.096 230.096230.096{\,}230.096 23. Thus, the optimal value of $r$ to maximize the probability is $r=1.146$. (b) Plot of the efficiency and probabilities of the heralded single photon as functions of $r$. The thick solid red and thin solid blue curves represent $E(r;-)$ and $E(r)$, respectively. The thick dashed red and thin dashed blue curves represent $P(1,1;r;-)$ and $P(1,1;r)$, respectively.
Figure 4: (a) Plot of ${\cal P}_{0}(\tilde{\tau},r,\alpha)$ as a function of $\tilde{\tau}$ and $r$ with $\alpha=10$. (b) Plot of ${\cal P}_{1}(\tilde{\tau},r,\alpha)$ as a function of $\tilde{\tau}$ and $r$ with $\alpha=10$.
Next in Fig. 3(a), we plot the actual probability that we obtain the $|1\rangle$ photon Fock state conditioned that we observe a heralding single photon for the initial state $|r;-\rangle_{a}$. Examining Fig. 3(a), we note that ${\cal N}_{-}(r)P(1,1;r;-)/4$ reaches its maximum value of 0.096 230.096230.096{\,}230.096 23 when $r=1.146$. At that point, the efficiency is $E(1.146;-)=0.9488$. Figure 3(b) shows plots of the efficiencies and probabilities of the heralded single photon as a function of $r$. We observe that $E(r;-)$ is larger than or equal to $E(r)$ for any $r$. Thus, as the heralded single-photon source, the superposition of oppositely squeezed states $|r;-\rangle$ is preferable to the typical squeezed state. It is useful to mention that the squeezed state $|r\rangle$ is a superposition of $\{|2n\rangle:n=0,1,2,...\}$ while contrastingly, the superpositions of the oppositely squeezed states $|r;+\rangle$ and $|r;-\rangle$ are superposition of $\{|2m\rangle:m=0,2,4,...\}$ and $\{|2m\rangle:m=1,3,5,...\}$, respectively. The Fock states present in these superpositions of squeezed states are further separated in photon number space than the usual squeezed state. If we can prepare $|r;-\rangle\propto c_{2}|2\rangle+c_{6}|6\rangle+...$ and inject that state into the 50-50 beam splitter, a transformation $|2\rangle_{a}|0\rangle_{b}\to(1/2)(|2\rangle_{a}|0\rangle_{b}+|0\rangle_{a}|2\rangle_{b})-(1/\sqrt{2})|1\rangle_{a}|1\rangle_{b}$ occurs and we obtain a state $|1\rangle_{a}|1\rangle_{b}$ with a maximal probability of $1/2$. Hence, we can implement the heralded single-photon source with higher efficiency. Figures 4(a) and (b) show 3D plots of the probabilities of generating $|r;-\rangle$ and a heralded single photon as functions of $\tilde{\tau}=2\kappa\tau$ and $r$, respectively, with fixing $\alpha$ at a specific value. In Fig. 4(a), we plot ${\cal P}_{0}$, the probability of generating $|r;-\rangle$, where
${\cal P}_{0}(\tilde{\tau},r,\alpha)=|_{12}\langle\Psi_{\mbox{\scriptsize out}}|r;-\rangle_{1}|-\alpha\rangle_{2}|^{2}.$ (12)
In Fig. 4(b), we plot ${\cal P}_{1}$, the probability of generating a heralded single photon, where
${\cal P}_{1}(\tilde{\tau},r,\alpha)=P(1,1,r;-){\cal P}_{0}(\tilde{\tau},r,\alpha).$ (13)
Looking at both of the plots, we observe sharp peaks at $\tilde{\tau}=\pi$.
II.3 Imperfections and errors
In any realistic implementation, there will be imperfections and so let us explore effects associated with the cross-phase modulation and also single-photon heralding detection. We have previously detailed the generation of $|r;-\rangle$ from the squeezed state $|r\rangle$ using the nonlinear cross-Kerr medium. For this generation, we had set duration $\tau$ for the nonlinear interaction as $2\kappa\tau=\pi$. Here, we consider a case where an inaccuracy occurs and such timing is not perfect and a phase error could be induced as $2\kappa\tau=\pi+\Delta\theta$. In this case
$|\Psi^{\prime}_{\mbox{\scriptsize out}}\rangle_{12}$ $=$ $\frac{1}{\sqrt{\cosh r}}\sum_{n=0}^{\infty}\frac{\sqrt{(2n)!}}{2^{n}n!}(-\tanh r)^{n}$ (14)
×$\times|2n\rangle_{1}|(-1)^{n}\alpha e^{-in\Delta\theta}\rangle_{2},$
where we assume $\alpha$ is real and greater than zero. To realize our heralded single-photon source, we want to generate a state of the form
$|\Psi^{(-)}_{\mbox{\scriptsize out}}\rangle_{12}=\frac{1}{2}{\cal N}_{-}(r)^{1/2}|r;-\rangle_{1}|-\alpha\rangle_{2},$ (15)
and so evaluate the effects caused by $\Delta\theta$, we define the following ratio that implies the generation rate of $|r;-\rangle$:
$R(r,\alpha,\Delta\theta)=\frac{|_{12}\langle\Psi^{(-)}_{\mbox{\scriptsize out}}|\Psi^{\prime}_{\mbox{\scriptsize out}}\rangle_{12}|^{2}}{\left.|_{12}\langle\Psi^{(-)}_{\mbox{\scriptsize out}}|\Psi^{\prime}_{\mbox{\scriptsize out}}\rangle_{12}|^{2}\right|_{\Delta\theta=0}}.$ (16)
Here $\Delta\theta$ is not a constant value but a probabilistic random variable that obeys Gaussian distribution whose mean value and standard deviation are given by zero and $\sigma$, respectively. Taking the average of $R(r,\alpha,\Delta\theta)$ with varying $\Delta\theta$ for many samples, we obtain $R(r,\alpha,\sigma)$. In Fig. 5(a), we plot $R(r,\alpha,\sigma)$ for $\alpha=9$, $10$, and $11$ with $r=0.725$. In the range of $0\leq\sigma\leq 0.001$, the function $\exp(-\lambda\sigma^{2})$ approximate to the curves of $R(r,\alpha,\sigma)$ well. Looking at those plots, we note that the curve of $R(r,\alpha,\sigma)$ gets closer to a horizontal as $\alpha$ becomes smaller. Thus, we become aware that the error of the inaccuracy of the duration is serious for large $\alpha$. Because the homodyne measurement requires a strong coherent state, we need to increase $|\alpha|$ for precise measurements. However, large $|\alpha|$ makes the single-photon emission vulnerable to the inaccuracy of duration $\tau$ and a trade-off emerges. Figure 5(b) shows a 3D plot of $R(r,\alpha,\sigma)$ with $\alpha=10$ as a function of $r$ and $\sigma$ where we note that, in the range where $r$ is larger than $1/2$, $R(r,10,\sigma)$ decreases as $\sigma$ gets larger. Fluctuations of the phase error are problematic when we try to obtain strong squeezing. If we want to generate a strongly squeezed state, we must ensure the accuracy of the cross-Kerr phase modulation giving us another trade-off.
Figure 5: (a) Plot of $R(r,\alpha,\sigma)$ as a function of $\sigma$ for $r=0.725$ and $\alpha=9$ (thick solid red line), $10$ (thick dashed blue line), and $11$ (thick dotted purple lines), respectively. The thin solid red, thin dashed blue, and thin dotted purple curves represent functions $\exp(-\lambda\sigma^{2})$ for $\lambda=3401\pm 3$, $5102\pm 2$, and $7360\pm 4$, respectively, that approximate the curves of $R(r,\alpha,\sigma)$ well for small $r$. (b) 3D plot of $R(r,\alpha,\sigma)$ as a function of both $r$ and $\sigma$ for $\alpha=10$. In the regime where $r$ is larger than $1/2$, $R(r,\alpha,\sigma)$ decreases as $\sigma$ gets larger.
Next, we consider inefficient single-photon detection. So far, we have assumed that we can use an ideal single-photon detector for the heralding measurement. The implementation of the single-photon detector is a cutting-edge technology with detection efficiencies above 90% regularly reported [29, 30].
We can model a “click/no click” detector of efficiency $\eta$ by the POVM $\{\hat{M},\hat{\mbox{\boldmath$I$}}-\hat{M}\}$ [3, 31, 32], where
$\hat{M}=\eta\sum_{k=1}^{\infty}(1-\eta)^{k-1}|k\rangle\langle k|,$ (17)
[…]
$Q(r,\eta;-)=\eta\sum_{n_{a}=1}^{\infty}\sum_{m_{b}=0}^{\infty}(1-\eta)^{n-1}P(n_{a},m_{b};r;-).$ (18)
The probability that the detector “clicks” and the heralded single photon is actually emitted is given by
$Q_{1}(r,\eta;-)=\eta\sum_{n_{a}=1}^{\infty}(1-\eta)^{n-1}P(n_{a},1;r;-).$ (19)
Then, we can compute the efficiency as
$E(r,\eta;-)=\frac{Q_{1}(r,\eta;-)}{Q(r,\eta;-)}.$ (20)
Figure 6: (a) Plot of $Q_{1}(r,\eta;-)$, $\tilde{Q}_{1}(r,\eta)$, and ${\cal N}_{-}(r)Q_{1}(r,\eta;-)/4$ as a function of $r$ for $\eta=0.9$, where they are represented by the thick solid red, thin solid blue, and thick dashed red curves, respectively. (b) Plot of $E(r,\eta;-)$ and $\tilde{E}(r,\eta)$ as a function of $r$ with $\eta=0.9$, where they are represented by the thick red and thin blue curves, respectively.
Figure 7: (a) 3D plots of $Q_{1}(r,\eta;-)$, $\tilde{Q}_{1}(r,\eta)$, and ${\cal N}_{-}(r)Q_{1}(r,\eta;-)/4$ as functions of $r$ and $\eta$, where they are represented by the curved red, blue, and green transparent surfaces, respectively. (b) 3D plots of $E(r,\eta;-)$ and $\tilde{E}(r,\eta)$ as functions of $r$ and $\eta$, where they are represented by the curved red and blue transparent surfaces, respectively. In the regime $r\leq 2$ and $0.7\leq\eta\leq 1$ we have $Q_{1}(r,\eta;-)>\tilde{Q}_{1}(r,\eta)>{\cal N}_{-}(r)Q_{1}(r,\eta;-)/4$ and $E(r,\eta;-)<\tilde{E}(r,\eta)$.
In Fig. 6(a), we plot $Q_{1}(r,\eta;-)$ and ${\cal N}_{-}(r)Q_{1}(r,\eta;-)/4$ as functions of $r$ for $\eta=0.9$, where ${\cal N}_{-}(r)Q_{1}(r,\eta;-)/4$ is the probability that the heralded single photon is actually emitted. Although the single-photon detector is imperfect ($\eta=0.9$), $Q_{1}(r,\eta;-)$ is nearly equal to $0.5$ for the small squeezing parameter $r$. This is one of the advantages that our method has.
Now in Fig. 6(b), we plot $E(r,\eta;-)$ as a function of $r$ for $\eta=0.9$. This plot shows that $E(r,\eta;-)$ is nearly equal to $0.95$ at $r=0$ and decreases gradually as $r$ becomes larger. In Fig. 7(a), we plot $Q_{1}(r,\eta;-)$ and ${\cal N}_{-}(r)Q_{1}(r,\eta;-)/4$ as functions of $r$ and $\eta$, where ${\cal N}_{-}(r)Q_{1}(r,\eta;-)/4$ is the probability that the heralded single photon is actually emitted. In the limit of $r\to 0$, $Q_{1}(r,\eta;-)$ is finite and non-zero but ${\cal N}_{-}(r)Q_{1}(r,\eta;-)/4$ approaches zero. In Fig. 7(b), we plot $E(r,\eta;-)$ as a function of $r$ and $\eta$. As $r$ increases and $\eta$ decreases, $E(r,\eta;-)$ becomes smaller gradually.
III Comparison with a heralded single-photon source from the two-mode squeezed state
Let us now review the generation of a heralded single photon using the pure two-mode squeezed state for comparison with our proposed method. Injecting two squeezed states individually onto a 50-50 beam splitter allows us to create the two-mode squeezed state
$|r\rangle_{a}|r\rangle_{b}\to\hat{S}_{ab}(ir)|0\rangle_{a}|0\rangle_{b},$ (21)
where $r$ is real. The probability we detect $n_{a}$ photons and $n_{b}$ photons is given by
$\tilde{P}(n_{a},n_{b};r)=\left|\langle n_{a},n_{b}|\hat{S}_{ab}(ir)|0_{a},0_{b}\rangle\right|^{2},$ (22)
meaning that we can write the probability for $n_{a}=1$ as $\tilde{P}(1,n_{b};r)=(\tanh r/\cosh r)^{2}\delta_{n_{b},1}$. If we observe a single photon in the first output mode, we obtain a single photon in the second output mode deterministically.
Figure 8: Plot of $P(1,1;r;-)$, $\tilde{P}(1,1;r)$, and ${\cal N}_{-}(r)P(1,1;r;-)/4$ as functions of $r$ which are represented as the thick solid red, thin solid blue, and thick dashed red curves, respectively. The graph of $\tilde{P}(1,1;r)$ reaches the maximum value $1/4$ for $r=\mbox{arcsech}(1/\sqrt{2})=0.881$ and reaches the minimum value 00 for $r=0$.
In Fig. 8, we plot $P(1,1;r;-)$, $\tilde{P}(1,1;r)$, and ${\cal N}_{-}(r)P(1,1;r;-)/4$ as a function of $r$. When $r=\mbox{\rm arcsech}(1/\sqrt{2})=0.881$, $\tilde{P}(1,1;r)$ achieves its maximum value of $1/4$. In the limit of $r\to 0$, $P(1,1;r;-)$ and $\tilde{P}(1,1;r)$ approach to $1/2$ and 00, respectively. The fact that $\tilde{P}(1,1;r)$ becomes zero in the limit of $r\to 0$ is a major defect for generation of the heralded single photon from the pure two-mode squeezed state. Under a practical condition where the squeezing parameter $r$ is small, we cannot generate the single photon efficiently. Thus, on condition of detecting $|-\alpha\rangle_{2}$ and $|1\rangle_{a}$, the superposition of oppositely squeezed states $|r;-\rangle$ is preferable to the pure two-mode squeezed state $\hat{S}_{ab}(ir)|0\rangle_{a}|0\rangle_{b}$ as a source of the heralded single photons. However, the actual probability of the emission of the single photons for $|r;-\rangle$ converges to zero as $r\to 0$.
Now if we prepare a factory where we can produce many light beams of $|r;-\rangle$ by the cross-Kerr nonlinear medium and the homodyne measurements, we obtain heralded single photons efficiently for small $r$ because of $\lim_{r\to 0}P(1,1;r;-)=0.5$. Moreover, looking at Fig. 3(b), we note that the efficiency $E(r;-)$ is close to unity for small $r$. Hence, if we construct the heralded single-photon source with $|r;-\rangle$, the emission of the single photons is nearly deterministic. Contrastingly, if we use the pure two-mode squeezed state $\hat{S}_{ab}(ir)|0\rangle_{a}|0\rangle_{b}$ for the heralded single-photon source, we cannot obtain the emission of the single photons efficiently because $\lim_{r\to 0}\tilde{P}(1,1;r)=0$. In Fig. 8, we note that $\tilde{P}(1,1;r)<P(1,1;r;-)$ for $0\leq r\leq 2$. Thus, $|r;-\rangle_{a}$ is preferable to $\hat{S}_{ab}(ir)|0\rangle_{a}|0\rangle_{b}$. Although we always obtain the heralded single photon when we detect the heralding single photon for the two-mode squeezed state and its efficiency is exactly equal to unity, looking at Fig. 3(b), we observe that $E(r;-)$ is nearly and sufficiently equal to unity, as well. Thus, regarding the efficiencies, the superposition of oppositely squeezed states $|r;-\rangle$ is as useful as the pure two-mode squeezed state.
Next, we consider the case where the single-photon detector is imperfect, and its “click” operator is given by Eq. (17). In this case, the click probability is
$\tilde{Q}(r,\eta)=\frac{2\eta\tanh^{2}r}{2-\eta[1-\cosh(2r)]}.$ (23)
The probability that the detector observes the click operator $\hat{M}$ and the heralded single photon is actually emitted is given by
$\tilde{Q}_{1}(r,\eta)=\eta\frac{\tanh^{2}r}{\cosh^{2}r}.$ (24)
Then, we can compute the efficiency as
$\tilde{E}(r,\eta)=\frac{\tilde{Q}_{1}(r,\eta)}{\tilde{Q}(r,\eta)}=\eta+(1-\eta)\cosh^{-2}r.$ (25)
In Figs. 6(a) and (b), we plot $\tilde{Q}_{1}(r,\eta)$ and $\tilde{E}(r,\eta)$ as functions of $r$ for $\eta=0.9$ while in Figs. 7(a) and (b), we plot $\tilde{Q}_{1}(r,\eta)$ and $\tilde{E}(r,\eta)$ as functions of $r$ and $\eta$. Figure 6(a) shows that $\tilde{Q}_{1}(r,\eta)$ is nearly equal to zero for small $r$. This characteristic is in contrast to our proposed method because $Q_{1}(r,\eta;-)$ is larger than $0.4$ for small $r$. Comparing $E(r,\eta;-)$ and $\tilde{E}(r,\eta)$ in Fig. 6(b), we cannot find a large difference between them for small $r$. For this point, both the two methods are comparable. Examining Fig. 7(a), we note that $Q_{1}(r,\eta;-)$ is larger than $\tilde{Q}_{1}(r,\eta)$ for $0\leq r\leq 2$ and $0.7\leq\eta\leq 1$. Moreover, $\lim_{r\to 0}Q_{1}(r,\eta;-)>0$ and $\lim_{r\to 0}\tilde{Q}_{1}(r,\eta)=0$ hold. Thus, for most realistic cases, the superposition of oppositely squeezed states is more beneficial to the heralded single-photon emission than the pure two-mode squeezed state. Looking at Fig. 7(b), we note that $E(r,\eta;-)<\tilde{E}(r,\eta)$ for $0\leq r\leq 2$ and $0.7\leq\eta\leq 1$. However, for larger $r$ and smaller $\eta$ the difference is small.
IV The second-order intensity correlation functions
Now, we consider the second-order intensity correlation functions $g^{(2)}(0)$ for the heralded single-photon emitters realized with the superposition of oppositely squeezed states $|r;-\rangle$ and the pure two-mode squeezed state $\hat{S}_{ab}(ir)|0\rangle_{a}|0\rangle_{b}$. We can measure the quality of the single-photon source by
$g^{(2)}(0)=\frac{\langle n(n-1)\rangle}{\langle n\rangle^{2}}.$ (26)
When an on-demand identical photon gun is realized, we obtain $g^{(2)}(0)=0$. For our proposed method, we have to calculate $g^{(2)}(0)$ numerically. By contrast, for the pure two-mode squeezed state,
$g^{(2)}(0)=-3+\frac{2}{\eta}+\eta+(1-\eta)\cosh 2r.$ (27)
Figure 9: (a) Plot of $g^{(2)}(0)$ for $|r;-\rangle$ (solid red curve) and $\hat{S}_{ab}(ir)|0\rangle_{a}|0\rangle_{b}$ (dashed blue curve) as a function of $r$ with $\eta=0.9$. (b) Plot of $g^{(2)}(0)$ as a function of $r$ and $\eta$.
We plot $g^{(2)}(0)$ for $|r;-\rangle$ and $\hat{S}_{ab}(ir)|0\rangle_{a}|0\rangle_{b}$ in Fig. 9. We observe that $g^{(2)}(0)$ is always less for $|r;-\rangle$ in the shown region. In particular for $|r;-\rangle$, we notice that $g^{(2)}(0)$ approaches zero as $r\to 0$ while for the two-mode squeezed state it is always greater than zero for $r\geq 0$. Thus, for $\eta=0.9$ and $0\leq r\leq 0.504$, the single-photon emitter realized by $|r;-\rangle$ has a better quality than the two-mode squeezed state. We can understand the reason why this difference emerges as follows. The state $|r;-\rangle$ is a superposition of photon-number states, $|2\rangle$, $|6\rangle$, $|10\rangle$, …. Thus, dividing the state $|r;-\rangle$ on two ports of a 50-50 beam splitter, the probability that the imperfect photon detector detects the state $|2\rangle_{a}$ is supressed. Thus, $g^{(2)}(0)$ hardly suffers from an imperfection caused by the term $\eta(1-\eta)|2\rangle_{a}{}_{a}\langle 2|$ included in the observable $\hat{M}$ given by Eq. (17). By contrast, the pure two-mode squeezed state is a superposition of the states, $|0\rangle_{a}|0\rangle_{b}$, $|1\rangle_{a}|1\rangle_{b}$, $|2\rangle_{a}|2\rangle_{b}$, …, so that it is vulnerable to the imperfect term $\eta(1-\eta)|2\rangle_{a}{}_{a}\langle 2|$. In Fig. 9(b), we plot $g^{(2)}(0)$ as a function of $r$ and $\eta$ for $0\leq r\leq 2$ and $0.7\leq\eta\leq 1$. We observe that $g^{(2)}(0)$ for our odd superposition of squeezed states is always smaller than that associated with the two-mode squeezed state for small $r$. This fact implies that the heralded single-photon emitter realized by $|r;-\rangle$ is more robust than that realized by $\hat{S}_{ab}(ir)|0\rangle_{a}|0\rangle_{b}$ under the condition that we must use the imperfect single-photon detector.
V Conclusion
In this paper, we proposed a new method to generate a heralded single-photon source created from a superposition of oppositely squeezed states. We outlined the method for producing this state by injecting squeezed and coherent light beams onto a cross-Kerr nonlinear medium and estimated the probability and efficiency of the emission of the heralded single photons. To evaluate the quality of our proposed method, we used the pure two-mode squeezed state for comparison. When the single-photon detector detects the heralding signals inaccurately, our method shows better performances for the second-order intensity correlation function $g^{(2)}(0)$ than the method of the pure two-mode squeezed states. This is because $|r\rangle-|-r\rangle$ is a superposition of photon-number states, $|2\rangle$, $|6\rangle$, $|10\rangle$, … and the probability amplitude of $|n\rangle_{a}|2\rangle_{b}$ for $n=0,1,2,...$ is suppressed when we inject it into the 50-50 beam splitter. That is an important advantage of scheme. However the generation of the odd superposition of squeezed states is more difficult leading to a natural trade-off.
Acknowledgment
This work was supported by MEXT Quantum Leap Flagship Program Grant No. JPMXS0120351339.

### v2

@@
-A heralded single-photon source based on superpositions of squeezed states
+Heralded single-photon source based on superpositions of squeezed states
 Abstract
 We propose a heralded single-photon source based on injecting a superposition of oppositely squeezed states onto a beam splitter.
-Our superposition of squeezed states is composed of only even photon number states (the number of photons is equal to $2,6,10,...$) meaning the probability for an emitted single photon given a heralded singe photon events is higher than what one can achieve from the usual two-mode squeezed state.
+Our superposition of squeezed states is composed of only even photon number states (the number of photons is equal to $2,6,10,...$) meaning the probability for an emitted single photon given as a heralded single-photon event is higher than what one can achieve from the usual two-mode squeezed state.
 This enables one to realize an enhanced heralded single-photon source.
-We discuss how to create this superposition of squeezed states utilizing single-mode squeezed states and the cross-Kerr nonlinearity.
-Our proposed method significantly improves the probability of emitting the heralded single photon compared to spontaneous parametric down conversion.
+We discuss how to create this superposition of squeezed states utilizing a single-mode squeezed state and the cross-Kerr nonlinearity.
+Our proposed method significantly improves the probability of emitting the heralded single photon compared to spontaneous parametric down-conversion.
 I Introduction
-Single photon sources are essential tools in many quantum information-based technologies including quantum key distribution protocols [1] and linear optical quantum computation [2, 3].
-As such there has been significant world-wide effort to propose and demonstrate on-demand single photon emitters.
-Quantum dots [4, 5, 6, 7, 8], nitrogen-vacancy centers in diamond [9, 10, 11], and negative silicon vacancy centers in diamond [12, 13, 14] have been proposed as solid state devices for realizing such emitter.
-Further cavity QED systems have been considered as a triggered single-photon source [15, 16, 17, 18] but the current work horse is still spontaneous parametric down conversion (SPDC) where we trigger off the idler photon [19, 20, 21].
+Single-photon sources are essential tools in many quantum information-based technologies including quantum key distribution protocols [1] and linear optical quantum computation [2, 3].
+As such there has been significant world-wide effort to propose and demonstrate on-demand single-photon emitters.
+Quantum dots [4, 5, 6, 7, 8], nitrogen-vacancy centers in diamond [9, 10, 11], and negative silicon vacancy centers in diamond [12, 13, 14] have been proposed as solid state devices for realizing such emitters.
+Further cavity QED systems have been considered as a triggered single-photon source [15, 16, 17, 18], but the current work horse is still spontaneous parametric down-conversion (SPDC) where we trigger off the idler photon [19, 20, 21, 22, 23, 24].
 However, the photon pair emission rate is low [$4\times 10^{-6}$ per pump photon using a periodically poled lithium niobate (PPLN) waveguide [25]].
-The essence of the heralded single-photon source is the generation of pairs of photons and the reason why we use SPDC.
+Implementation of heralded single-photon sources with the SPDC is state-of-the-art technology and is being studied with great effort currently [26].
+A review of the heralded single-photon sources was given in [27, 28].
+Improvement of the heralded single-photon source with photon-number-resolving detectors was examined in Refs.
+[29, 30].
+Suppression of multiphoton events in a cavity-enhanced SPDC via the photon-blockade effect for the heralded single-photon source has been proposed [31], and a technique of heralding based on the detection of zero photons was investigated [32].
+The essence of the heralded single-photon source is the generation of pairs of photons and is the reason we use SPDC.
 A more efficient method for generating entangled photons is the injection of squeezed light into a 50-50 beam splitter [33].
 The resulting two-mode squeezed state source can in principle generate pairs of photons with a much higher success probability.
-There is however still room for improvement.
-In this paper, we introduce a new and more efficient approach based on superposition of squeezed state.
+Using an imperfect single-photon detector whose efficiency is equal to $0.9$, the pure two-mode squeezed state with a squeezing parameter $0.5$ gives the photon pair emission rate $0.151$.
+There is, however, still room for improvement.
+In this paper, we introduce a more efficient approach based on superpositions of squeezed states.
 We investigate how to generate a pair of photons by injecting an odd superposition of squeezed states onto a 50-50 beam splitter [34].
-In particular we show that we can create this superposition using a single-mode squeezed state source and a cross-Kerr nonlinear medium although its heralded success probability is less than $1/2$ [25].
+In particular we show that we can create this superposition using a single-mode squeezed state source and a cross-Kerr nonlinear medium, although its heralded success probability is less than $1/2$ [35, 36].
 Finally, we determine the second-order intensity correlation functions $g^{(2)}(0)$ for the single-photon emissions of the superposition of oppositely squeezed states and the two-mode squeezed state.
@@
 II, we begin by discussing the generation of a pair of photons using our superposition of oppositely squeezed states determining in Secs.
-II.1 and II.2 the probability and efficiency of emission of the heralded single photon source.
-Then Sec.
+II.1 and II.2 the probability and conditional probability of emission of the heralded single-photon source.
+Then in Sec.
 II.3, we consider errors induced by imperfect operations of devices.
@@
 III, we compare our proposed heralded source with that from the pure two-mode squeezed state.
-Now in Sec.
+In Sec.
 IV, we determine the second-order intensity correlation functions for our source using imperfect single-photon detectors.
@@
 In our scheme, we assume that we have a stable source of squeezed states $|r\rangle$, as the generation of squeezed states is a mature technology.
-A typical squeezing level of squeezed light used for experiments is given by $6.3$ dB which corresponds to the squeezing parameter $|r|=0.725$ [27].
+A typical squeezing level of squeezed light used for experiments is given by $6.3$ dB, which corresponds to the squeezing parameter $|r|=0.725$ [38].
 In the current paper, we typically use this parameter for showing examples but note that the detection of a $15$ dB squeezed state of light was demonstrated experimentally [39].
-Figure 1: Schematic illustration of a heralded single photon source using squeezed state and a cross-Kerr medium.
-II.1 Heralded single-photon probability and efficiency
+In the proposed method, it is critical to adjust the duration of the nonlinear interaction time $\tau$.
+To realize a precise adjustment of $\tau$, we can fabricate a cavity of the cross-Kerr nonlinear medium, which is caught between two half-silvered mirrors.
+We can vary $\tau$ by changing the transmittance of mirrors.
+To obtain the phase factor $e^{-i2n\kappa\tau}$with $2\kappa\tau=\pi$ for Eq.
+(3) precisely, $\kappa$ must be large enough.
+Figure 1: Schematic illustration of a heralded single-photon source using a squeezed state and a cross-Kerr medium.
+II.1 Heralded single-photon probability and conditional probability
 Now let us establish the probability of successfully emitting our heralded single photon.
@@
 $P(n_{a},n_{b};r)=|\langle n_{a},n_{b}|\hat{S}_{ab}(-\frac{r}{2})\hat{S}_{a}(\frac{r}{2})\hat{S}_{b}(\frac{r}{2})|0_{a},0_{b}\rangle|^{2}.$ (9)
-We can then define our generation efficiencies $E(r;\pm)$ and $E(r)$ as
-$E(r;\pm)=P(1,1;r;\pm)/\sum_{n=0}^{\infty}P(1,n;r;\pm).$ (10)
-$E(r)=P(1,1;r)/\sum_{n=0}^{\infty}P(1,n;r).$ (11)
-If we observe |±$|\pm\alpha\rangle_{2}$ with the homodyne measurement, inject $|r;\pm\rangle_{a}$ into the beam splitter, and detect the heralding single photon in one port, we then obtain a single photon from the other port with probability $E(r;\pm)$.
-Similarly if we inject $|r\rangle_{a}$ into the beam splitter and detect the heralding single photon, we obtain a single photon with probability $E(r)$.
+We can then define our conditional probabilities $P_{\mbox{\scriptsize c}}(r;\pm)$ and $P_{\mbox{\scriptsize c}}(r)$ as
+$P_{\mbox{\scriptsize c}}(r;\pm)=P(1,1;r;\pm)/\sum_{n=0}^{\infty}P(1,n;r;\pm).$ (10)
+$P_{\mbox{\scriptsize c}}(r)=P(1,1;r)/\sum_{n=0}^{\infty}P(1,n;r).$ (11)
+If we observe |±$|\pm\alpha\rangle_{2}$ with the homodyne measurement, inject $|r;\pm\rangle_{a}$ into the beam splitter, and detect the heralding single photon in one port, we then obtain a single photon from the other port with probabilities $P_{\mbox{\scriptsize c}}(r;\pm)$.
+Thus, we can regard $P_{\mbox{\scriptsize c}}(r;\pm)$ as the conditional probabilities of single-photon emission after the postselection with the measurement outcomes on the other port.
+Similarly if we inject $|r\rangle_{a}$ into the beam splitter and detect the heralding single photon, we obtain a single photon with conditional probability $P_{\mbox{\scriptsize c}}(r)$.
 II.2 Ideal single photon detection
@@
 2, we note that $P(1,1;r;-)=0.453$, $P(1,5;r;-)=7.85\times 10^{-3}$, and $P(1,n;r;-)=0$ for $n=0,2,3,4$.
-Moreover, we obtain $E(r;-)=0.983$.
+Moreover, we obtain $P_{\mbox{\scriptsize c}}(r;-)=0.983$.
 If we generate the superposition of the oppositely squeezed states $|r;-\rangle_{a}$, it should allow us to realize the heralded single-photon source with a high probability and so be a good alternative to the SPDC.
-By contrast, if we prepare the squeezed state $|r\rangle_{a}$, we obtain $P(1,1;r)=7.54\times 10^{-2}$ with $E(r)=0.859$.
+By contrast, if we prepare the squeezed state $|r\rangle_{a}$, we obtain $P(1,1;r)=7.54\times 10^{-2}$ with $P_{\mbox{\scriptsize c}}(r)=0.859$.
 Looking at the plot of $|r;-\rangle$, the following question comes to mind.
@@
 3.
-Figure 2: Plot of the probability $P(1,n;r)$ and $P(1,n;r;\pm)$ we generate an $n$ photon state $|n\rangle$ in one mode and a heralding photon in the other for the injected states $|r\rangle_{a}$ and $|r;\pm\rangle_{a}$ with $r=0.725$.
+Figure 2: Plot of the probability $P(1,n;r)$ and $P(1,n;r;\pm)$.
+We generate an $n$ photon state $|n\rangle$ in one mode and a heralding photon in the other for the injected states $|r\rangle_{a}$ and $|r;\pm\rangle_{a}$ with $r=0.725$.
 The dotted blue, solid red, and dashed purple line graphs represent $P(1,n;r)$, $P(1,n;r;-)$, and $P(1,n;r;+)$, respectively.
 The line graph of $P(1,n;r;+)$ overlaps the horizontal axis of $n$ mostly.
-Figure 3: (a) Plot of the probability ${\cal N}_{-}(r)P(1,1;r;-)/4$ we obtain a heralded single photon by injecting $|r;-\rangle_{a}$ for $0\leq r\leq 2$.
+Figure 3: (a) Plot of the probability ${\cal N}_{-}(r)P(1,1;r;-)/4$.
+We obtain a heralded single photon by injecting $|r;-\rangle_{a}$ for $0\leq r\leq 2$.
 When $r=1.146$, the graph reaches its maximum value 0.096 230.096230.096{\,}230.096 23.
 Thus, the optimal value of $r$ to maximize the probability is $r=1.146$.
-(b) Plot of the efficiency and probabilities of the heralded single photon as functions of $r$.
-The thick solid red and thin solid blue curves represent $E(r;-)$ and $E(r)$, respectively.
+(b) Plot of the probabilities of the heralded single photon as functions of $r$.
+The thick solid red and thin solid blue curves represent $P_{\mbox{\scriptsize c}}(r;-)$ and $P_{\mbox{\scriptsize c}}(r)$, respectively.
 The thick dashed red and thin dashed blue curves represent $P(1,1;r;-)$ and $P(1,1;r)$, respectively.
@@
 3(a), we note that ${\cal N}_{-}(r)P(1,1;r;-)/4$ reaches its maximum value of 0.096 230.096230.096{\,}230.096 23 when $r=1.146$.
-At that point, the efficiency is $E(1.146;-)=0.9488$.
-Figure 3(b) shows plots of the efficiencies and probabilities of the heralded single photon as a function of $r$.
-We observe that $E(r;-)$ is larger than or equal to $E(r)$ for any $r$.
+At that point, the conditional probability is $P_{\mbox{\scriptsize c}}(1.146;-)=0.9488$.
+Figure 3(b) shows plots of the probabilities of the heralded single photon as a function of $r$.
+We observe that $P_{\mbox{\scriptsize c}}(r;-)$ is larger than or equal to $P_{\mbox{\scriptsize c}}(r)$ for any $r$.
 Thus, as the heralded single-photon source, the superposition of oppositely squeezed states $|r;-\rangle$ is preferable to the typical squeezed state.
-It is useful to mention that the squeezed state $|r\rangle$ is a superposition of $\{|2n\rangle:n=0,1,2,...\}$ while contrastingly, the superpositions of the oppositely squeezed states $|r;+\rangle$ and $|r;-\rangle$ are superposition of $\{|2m\rangle:m=0,2,4,...\}$ and $\{|2m\rangle:m=1,3,5,...\}$, respectively.
+It is useful to mention that the squeezed state $|r\rangle$ is a superposition of $\{|2n\rangle:n=0,1,2,...\}$ while contrastingly, the superpositions of the oppositely squeezed states $|r;+\rangle$ and $|r;-\rangle$ are superpositions of $\{|2m\rangle:m=0,2,4,...\}$ and $\{|2m\rangle:m=1,3,5,...\}$, respectively.
 The Fock states present in these superpositions of squeezed states are further separated in photon number space than the usual squeezed state.
-If we can prepare $|r;-\rangle\propto c_{2}|2\rangle+c_{6}|6\rangle+...$ and inject that state into the 50-50 beam splitter, a transformation $|2\rangle_{a}|0\rangle_{b}\to(1/2)(|2\rangle_{a}|0\rangle_{b}+|0\rangle_{a}|2\rangle_{b})-(1/\sqrt{2})|1\rangle_{a}|1\rangle_{b}$ occurs and we obtain a state $|1\rangle_{a}|1\rangle_{b}$ with a maximal probability of $1/2$.
-Hence, we can implement the heralded single-photon source with higher efficiency.
-Figures 4(a) and (b) show 3D plots of the probabilities of generating $|r;-\rangle$ and a heralded single photon as functions of $\tilde{\tau}=2\kappa\tau$ and $r$, respectively, with fixing $\alpha$ at a specific value.
+If we can prepare $|r;-\rangle\propto c_{2}|2\rangle+c_{6}|6\rangle+...$ and inject that state into the 50-50 beam splitter, a transformation $|2\rangle_{a}|0\rangle_{b}\to(1/2)(|2\rangle_{a}|0\rangle_{b}+|0\rangle_{a}|2\rangle_{b})-(1/\sqrt{2})|1\rangle_{a}|1\rangle_{b}$ occurs, and we obtain a state $|1\rangle_{a}|1\rangle_{b}$ with a maximal probability of $1/2$.
+Hence, we can implement the heralded single-photon source with higher conditional probability.
+Figures 4(a) and 4(b) show 3D plots of the probabilities of generating $|r;-\rangle$ and a heralded single photon as functions of $\tilde{\tau}=2\kappa\tau$ and $r$, respectively, with fixing $\alpha$ at a specific value.
 In Fig.
@@
 ${\cal P}_{1}(\tilde{\tau},r,\alpha)=P(1,1,r;-){\cal P}_{0}(\tilde{\tau},r,\alpha).$ (13)
-Looking at both of the plots, we observe sharp peaks at $\tilde{\tau}=\pi$.
+Looking at both plots, we observe sharp peaks at $\tilde{\tau}=\pi$.
 II.3 Imperfections and errors
-In any realistic implementation, there will be imperfections and so let us explore effects associated with the cross-phase modulation and also single-photon heralding detection.
+In any realistic implementation, there will be imperfections, and so let us explore effects associated with the cross-phase modulation and also single-photon heralding detection.
 We have previously detailed the generation of $|r;-\rangle$ from the squeezed state $|r\rangle$ using the nonlinear cross-Kerr medium.
 For this generation, we had set duration $\tau$ for the nonlinear interaction as $2\kappa\tau=\pi$.
-Here, we consider a case where an inaccuracy occurs and such timing is not perfect and a phase error could be induced as $2\kappa\tau=\pi+\Delta\theta$.
+Here, we consider a case where an inaccuracy occurs; such timing is not perfect and a phase error could be induced as $2\kappa\tau=\pi+\Delta\theta$.
 In this case
@@
 $|\Psi^{(-)}_{\mbox{\scriptsize out}}\rangle_{12}=\frac{1}{2}{\cal N}_{-}(r)^{1/2}|r;-\rangle_{1}|-\alpha\rangle_{2},$ (15)
-and so evaluate the effects caused by $\Delta\theta$, we define the following ratio that implies the generation rate of $|r;-\rangle$:
+and so to evaluate the effects caused by $\Delta\theta$, we define the following ratio that implies the decrease in generation probability of $|r;-\rangle$ due to $\Delta\theta$:
 $R(r,\alpha,\Delta\theta)=\frac{|_{12}\langle\Psi^{(-)}_{\mbox{\scriptsize out}}|\Psi^{\prime}_{\mbox{\scriptsize out}}\rangle_{12}|^{2}}{\left.|_{12}\langle\Psi^{(-)}_{\mbox{\scriptsize out}}|\Psi^{\prime}_{\mbox{\scriptsize out}}\rangle_{12}|^{2}\right|_{\Delta\theta=0}}.$ (16)
 Here $\Delta\theta$ is not a constant value but a probabilistic random variable that obeys Gaussian distribution whose mean value and standard deviation are given by zero and $\sigma$, respectively.
-Taking the average of $R(r,\alpha,\Delta\theta)$ with varying $\Delta\theta$ for many samples, we obtain $R(r,\alpha,\sigma)$.
+Taking the average of $R(r,\alpha,\Delta\theta)$ with varying $\Delta\theta$ for many samples, we numerically obtain
+$R(r,\alpha,\sigma)=\int_{-\infty}^{\infty}d(\Delta\theta)\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp(-\frac{\Delta\theta^{2}}{2\sigma^{2}})R(r,\alpha,\Delta\theta).$ (17)
 In Fig.
 5(a), we plot $R(r,\alpha,\sigma)$ for $\alpha=9$, $10$, and $11$ with $r=0.725$.
-In the range of $0\leq\sigma\leq 0.001$, the function $\exp(-\lambda\sigma^{2})$ approximate to the curves of $R(r,\alpha,\sigma)$ well.
+In the range of $0\leq\sigma\leq 0.001$, the function $\exp(-\lambda\sigma^{2})$ approximates to the curves of $R(r,\alpha,\sigma)$ well.
 Looking at those plots, we note that the curve of $R(r,\alpha,\sigma)$ gets closer to a horizontal as $\alpha$ becomes smaller.
@@
 Because the homodyne measurement requires a strong coherent state, we need to increase $|\alpha|$ for precise measurements.
-However, large $|\alpha|$ makes the single-photon emission vulnerable to the inaccuracy of duration $\tau$ and a trade-off emerges.
+However, large $|\alpha|$ makes the single-photon emission vulnerable to the inaccuracy of duration $\tau$, and a trade-off emerges.
 Figure 5(b) shows a 3D plot of $R(r,\alpha,\sigma)$ with $\alpha=10$ as a function of $r$ and $\sigma$ where we note that, in the range where $r$ is larger than $1/2$, $R(r,10,\sigma)$ decreases as $\sigma$ gets larger.
@@
 If we want to generate a strongly squeezed state, we must ensure the accuracy of the cross-Kerr phase modulation giving us another trade-off.
+In Fig.
+5(a), $R(r,\alpha,\sigma)$ departs from unity considerably for $\sigma=0.004$.
+This standard deviation corresponds to a fluctuation $\Delta\tau=\tau\sigma/\pi$.
+This evaluation suggests to us that we must fix $\tau$ with the accuracy $\Delta\tau/\tau\sim 10^{-3}$, and it is well within our reach experimentally.
 Figure 5: (a) Plot of $R(r,\alpha,\sigma)$ as a function of $\sigma$ for $r=0.725$ and $\alpha=9$ (thick solid red line), $10$ (thick dashed blue line), and $11$ (thick dotted purple lines), respectively.
@@
 […]
-$Q(r,\eta;-)=\eta\sum_{n_{a}=1}^{\infty}\sum_{m_{b}=0}^{\infty}(1-\eta)^{n-1}P(n_{a},m_{b};r;-).$ (18)
+$P_{\mbox{\scriptsize click}}(r,\eta;-)=\eta\sum_{n_{a}=1}^{\infty}\sum_{m_{b}=0}^{\infty}(1-\eta)^{n_{a}-1}P(n_{a},m_{b};r;-).$ (19)
 The probability that the detector “clicks” and the heralded single photon is actually emitted is given by
-$Q_{1}(r,\eta;-)=\eta\sum_{n_{a}=1}^{\infty}(1-\eta)^{n-1}P(n_{a},1;r;-).$ (19)
-Then, we can compute the efficiency as
-$E(r,\eta;-)=\frac{Q_{1}(r,\eta;-)}{Q(r,\eta;-)}.$ (20)
-Figure 6: (a) Plot of $Q_{1}(r,\eta;-)$, $\tilde{Q}_{1}(r,\eta)$, and ${\cal N}_{-}(r)Q_{1}(r,\eta;-)/4$ as a function of $r$ for $\eta=0.9$, where they are represented by the thick solid red, thin solid blue, and thick dashed red curves, respectively.
-(b) Plot of $E(r,\eta;-)$ and $\tilde{E}(r,\eta)$ as a function of $r$ with $\eta=0.9$, where they are represented by the thick red and thin blue curves, respectively.
-Figure 7: (a) 3D plots of $Q_{1}(r,\eta;-)$, $\tilde{Q}_{1}(r,\eta)$, and ${\cal N}_{-}(r)Q_{1}(r,\eta;-)/4$ as functions of $r$ and $\eta$, where they are represented by the curved red, blue, and green transparent surfaces, respectively.
-(b) 3D plots of $E(r,\eta;-)$ and $\tilde{E}(r,\eta)$ as functions of $r$ and $\eta$, where they are represented by the curved red and blue transparent surfaces, respectively.
-In the regime $r\leq 2$ and $0.7\leq\eta\leq 1$ we have $Q_{1}(r,\eta;-)>\tilde{Q}_{1}(r,\eta)>{\cal N}_{-}(r)Q_{1}(r,\eta;-)/4$ and $E(r,\eta;-)<\tilde{E}(r,\eta)$.
+$P_{\mbox{\scriptsize click},1}(r,\eta;-)=\eta\sum_{n_{a}=1}^{\infty}(1-\eta)^{n_{a}-1}P(n_{a},1;r;-).$ (20)
+Then, we can compute the conditional probability as
+$P_{\mbox{\scriptsize click,c}}(r,\eta;-)=\frac{P_{\mbox{\scriptsize click},1}(r,\eta;-)}{P_{\mbox{\scriptsize click}}(r,\eta;-)}.$ (21)
+Figure 6: (a) Plot of $P_{\mbox{\scriptsize click},1}(r,\eta;-)$, $\tilde{P}_{\mbox{\scriptsize click},1}(r,\eta)$ defined in Eq.
+(25), and ${\cal N}_{-}(r)P_{\mbox{\scriptsize click},1}(r,\eta;-)/4$ as a function of $r$ for $\eta=0.9$, where they are represented by the thick solid red, thin solid blue, and thick dashed red curves, respectively.
+(b) Plot of $P_{\mbox{\scriptsize click,c}}(r,\eta;-)$ and $\tilde{P}_{\mbox{\scriptsize click,c}}(r,\eta)$ defined in Eq.
+(26) as a function of $r$ with $\eta=0.9$, where they are represented by the thick red and thin blue curves, respectively.
+Figure 7: (a) 3D plots of $P_{\mbox{\scriptsize click},1}(r,\eta;-)$, $\tilde{P}_{\mbox{\scriptsize click},1}(r,\eta)$, and ${\cal N}_{-}(r)P_{\mbox{\scriptsize click},1}(r,\eta;-)/4$ as functions of $r$ and $\eta$, where they are represented by the curved red, blue, and green transparent surfaces, respectively.
+(b) 3D plots of $P_{\mbox{\scriptsize click,c}}(r,\eta;-)$ and $\tilde{P}_{\mbox{\scriptsize click,c}}(r,\eta)$ as functions of $r$ and $\eta$, where they are represented by the curved red and blue transparent surfaces, respectively.
+In the regime $r\leq 2$ and $0.7\leq\eta\leq 1$ we have $P_{\mbox{\scriptsize click},1}(r,\eta;-)>\tilde{P}_{\mbox{\scriptsize click},1}(r,\eta)>{\cal N}_{-}(r)P_{\mbox{\scriptsize click},1}(r,\eta;-)/4$ and $P_{\mbox{\scriptsize click,c}}(r,\eta;-)<\tilde{P}_{\mbox{\scriptsize click,c}}(r,\eta)$.
 In Fig.
-6(a), we plot $Q_{1}(r,\eta;-)$ and ${\cal N}_{-}(r)Q_{1}(r,\eta;-)/4$ as functions of $r$ for $\eta=0.9$, where ${\cal N}_{-}(r)Q_{1}(r,\eta;-)/4$ is the probability that the heralded single photon is actually emitted.
-Although the single-photon detector is imperfect ($\eta=0.9$), $Q_{1}(r,\eta;-)$ is nearly equal to $0.5$ for the small squeezing parameter $r$.
+6(a), we plot $P_{\mbox{\scriptsize click},1}(r,\eta;-)$ and ${\cal N}_{-}(r)P_{\mbox{\scriptsize click},1}(r,\eta;-)/4$ as functions of $r$ for $\eta=0.9$, where ${\cal N}_{-}(r)P_{\mbox{\scriptsize click},1}(r,\eta;-)/4$ is the probability that the heralded single photon is actually emitted.
+Although the single-photon detector is imperfect ($\eta=0.9$), $P_{\mbox{\scriptsize click},1}(r,\eta;-)$ is nearly equal to $0.5$ for the small squeezing parameter $r$.
 This is one of the advantages that our method has.
-Now in Fig.
-6(b), we plot $E(r,\eta;-)$ as a function of $r$ for $\eta=0.9$.
-This plot shows that $E(r,\eta;-)$ is nearly equal to $0.95$ at $r=0$ and decreases gradually as $r$ becomes larger.
 In Fig.
-7(a), we plot $Q_{1}(r,\eta;-)$ and ${\cal N}_{-}(r)Q_{1}(r,\eta;-)/4$ as functions of $r$ and $\eta$, where ${\cal N}_{-}(r)Q_{1}(r,\eta;-)/4$ is the probability that the heralded single photon is actually emitted.
-In the limit of $r\to 0$, $Q_{1}(r,\eta;-)$ is finite and non-zero but ${\cal N}_{-}(r)Q_{1}(r,\eta;-)/4$ approaches zero.
+6(b), we plot $P_{\mbox{\scriptsize click,c}}(r,\eta;-)$ as a function of $r$ for $\eta=0.9$.
+This plot shows that $P_{\mbox{\scriptsize click,c}}(r,\eta;-)$ is nearly equal to $0.95$ at $r=0$ and decreases gradually as $r$ becomes larger.
 In Fig.
-7(b), we plot $E(r,\eta;-)$ as a function of $r$ and $\eta$.
-As $r$ increases and $\eta$ decreases, $E(r,\eta;-)$ becomes smaller gradually.
+7(a), we plot $P_{\mbox{\scriptsize click},1}(r,\eta;-)$ and ${\cal N}_{-}(r)P_{\mbox{\scriptsize click},1}(r,\eta;-)/4$ as functions of $r$ and $\eta$, where ${\cal N}_{-}(r)P_{\mbox{\scriptsize click},1}(r,\eta;-)/4$ is the probability that the heralded single photon is actually emitted.
+In the limit of $r\to 0$, $P_{\mbox{\scriptsize click},1}(r,\eta;-)$ is finite and nonzero, but ${\cal N}_{-}(r)P_{\mbox{\scriptsize click},1}(r,\eta;-)/4$ approaches zero.
+In Fig.
+7(b), we plot $P_{\mbox{\scriptsize click,c}}(r,\eta;-)$ as a function of $r$ and $\eta$.
+As $r$ increases and $\eta$ decreases, $P_{\mbox{\scriptsize click,c}}(r,\eta;-)$ becomes smaller gradually.
 III Comparison with a heralded single-photon source from the two-mode squeezed state
@@
 The graph of $\tilde{P}(1,1;r)$ reaches the maximum value $1/4$ for $r=\mbox{arcsech}(1/\sqrt{2})=0.881$ and reaches the minimum value 00 for $r=0$.
+We put a small frame of plots of ${\cal N}_{-}(r)P(1,1;r;-)/4$ and $\tilde{P}(1,1;r)$ for $0\leq r\leq 0.006$ in the main frame.
+These plots show that ${\cal N}_{-}(r)P(1,1;r;-)/4$ and $\tilde{P}(1,1;r)$ are larger than $4\times 10^{-6}$ for $r\geq 0.004$.
+Thus, the heralded single-photon sources of our scheme and the pure two-mode squeezed state are more efficient than the SPDC even for small squeezing parameter $r$.
 In Fig.
@@
 When $r=\mbox{\rm arcsech}(1/\sqrt{2})=0.881$, $\tilde{P}(1,1;r)$ achieves its maximum value of $1/4$.
-In the limit of $r\to 0$, $P(1,1;r;-)$ and $\tilde{P}(1,1;r)$ approach to $1/2$ and 00, respectively.
-The fact that $\tilde{P}(1,1;r)$ becomes zero in the limit of $r\to 0$ is a major defect for generation of the heralded single photon from the pure two-mode squeezed state.
-Under a practical condition where the squeezing parameter $r$ is small, we cannot generate the single photon efficiently.
-Thus, on condition of detecting $|-\alpha\rangle_{2}$ and $|1\rangle_{a}$, the superposition of oppositely squeezed states $|r;-\rangle$ is preferable to the pure two-mode squeezed state $\hat{S}_{ab}(ir)|0\rangle_{a}|0\rangle_{b}$ as a source of the heralded single photons.
-However, the actual probability of the emission of the single photons for $|r;-\rangle$ converges to zero as $r\to 0$.
+In the limit of $r\to 0$, $P(1,1;r;-)$ and $\tilde{P}(1,1;r)$ approach $1/2$ and 00, respectively.
+In the limit of $r\to 0$, both ${\cal N}_{-}(r)P(1,1;r;-)/4$ and $\tilde{P}(1,1;r)$ approach zero, sharing the same defect.
+However, in a practical circumstance where the squeezing parameter $r$ is small, our scheme with the postselection of measuring $|-\alpha\rangle_{2}$ shown in Fig.
+1 gives the probability $P(1,1;r;-)$ to be nearly equal to $1/2$ around $r=0$.
+Hence, the superposition of oppositely squeezed states $|r;-\rangle$ is preferable to the pure two-mode squeezed state $\hat{S}_{ab}(ir)|0\rangle_{a}|0\rangle_{b}$ as a source of the heralded single photons.
 Now if we prepare a factory where we can produce many light beams of $|r;-\rangle$ by the cross-Kerr nonlinear medium and the homodyne measurements, we obtain heralded single photons efficiently for small $r$ because of $\lim_{r\to 0}P(1,1;r;-)=0.5$.
 Moreover, looking at Fig.
-3(b), we note that the efficiency $E(r;-)$ is close to unity for small $r$.
+3(b), we note that the conditional probability $P_{\mbox{\scriptsize c}}(r;-)$ is close to unity for small $r$.
 Hence, if we construct the heralded single-photon source with $|r;-\rangle$, the emission of the single photons is nearly deterministic.
@@
 Thus, $|r;-\rangle_{a}$ is preferable to $\hat{S}_{ab}(ir)|0\rangle_{a}|0\rangle_{b}$.
-Although we always obtain the heralded single photon when we detect the heralding single photon for the two-mode squeezed state and its efficiency is exactly equal to unity, looking at Fig.
-3(b), we observe that $E(r;-)$ is nearly and sufficiently equal to unity, as well.
-Thus, regarding the efficiencies, the superposition of oppositely squeezed states $|r;-\rangle$ is as useful as the pure two-mode squeezed state.
+Although we always obtain the heralded single photon when we detect the heralding single photon for the two-mode squeezed state and its conditional probability is exactly equal to unity, looking at Fig.
+3(b), we observe that $P_{\mbox{\scriptsize c}}(r;-)$ is nearly and sufficiently equal to unity, as well.
+Thus, regarding the conditional probabilities, the superposition of oppositely squeezed states $|r;-\rangle$ is as useful as the pure two-mode squeezed state.
 Next, we consider the case where the single-photon detector is imperfect, and its “click” operator is given by Eq. (18). In this case, the click probability is
-$\tilde{Q}(r,\eta)=\frac{2\eta\tanh^{2}r}{2-\eta[1-\cosh(2r)]}.$ (23)
+$\tilde{P}_{\mbox{\scriptsize click}}(r,\eta)=\frac{2\eta\tanh^{2}r}{2-\eta[1-\cosh(2r)]}.$ (24)
 The probability that the detector observes the click operator $\hat{M}$ and the heralded single photon is actually emitted is given by
-$\tilde{Q}_{1}(r,\eta)=\eta\frac{\tanh^{2}r}{\cosh^{2}r}.$ (24)
-Then, we can compute the efficiency as
-$\tilde{E}(r,\eta)=\frac{\tilde{Q}_{1}(r,\eta)}{\tilde{Q}(r,\eta)}=\eta+(1-\eta)\cosh^{-2}r.$ (25)
+$\tilde{P}_{\mbox{\scriptsize click},1}(r,\eta)=\eta\frac{\tanh^{2}r}{\cosh^{2}r}.$ (25)
+Then, we can compute the conditional probability as
+$\tilde{P}_{\mbox{\scriptsize click,c}}(r,\eta)=\frac{\tilde{P}_{\mbox{\scriptsize click},1}(r,\eta)}{\tilde{P}_{\mbox{\scriptsize click}}(r,\eta)}=\eta+(1-\eta)\cosh^{-2}r.$ (26)
 In Figs.
-6(a) and (b), we plot $\tilde{Q}_{1}(r,\eta)$ and $\tilde{E}(r,\eta)$ as functions of $r$ for $\eta=0.9$ while in Figs.
-7(a) and (b), we plot $\tilde{Q}_{1}(r,\eta)$ and $\tilde{E}(r,\eta)$ as functions of $r$ and $\eta$.
-Figure 6(a) shows that $\tilde{Q}_{1}(r,\eta)$ is nearly equal to zero for small $r$.
-This characteristic is in contrast to our proposed method because $Q_{1}(r,\eta;-)$ is larger than $0.4$ for small $r$.
-Comparing $E(r,\eta;-)$ and $\tilde{E}(r,\eta)$ in Fig.
+6(a) and 6(b), we plot $\tilde{P}_{\mbox{\scriptsize click},1}(r,\eta)$ and $\tilde{P}_{\mbox{\scriptsize click,c}}(r,\eta)$ as functions of $r$ for $\eta=0.9$ while in Figs.
+7(a) and 7(b), we plot $\tilde{P}_{\mbox{\scriptsize click},1}(r,\eta)$ and $\tilde{P}_{\mbox{\scriptsize click,c}}(r,\eta)$ as functions of $r$ and $\eta$.
+Figure 6(a) shows that $\tilde{P}_{\mbox{\scriptsize click},1}(r,\eta)$ is nearly equal to zero for small $r$.
+This characteristic is in contrast to our proposed method because $P_{\mbox{\scriptsize click},1}(r,\eta;-)$ is larger than $0.4$ for small $r$.
+Comparing $P_{\mbox{\scriptsize click,c}}(r,\eta;-)$ and $\tilde{P}_{\mbox{\scriptsize click,c}}(r,\eta)$ in Fig.
 6(b), we cannot find a large difference between them for small $r$.
-For this point, both the two methods are comparable.
+For this point, both methods are comparable.
 Examining Fig.
-7(a), we note that $Q_{1}(r,\eta;-)$ is larger than $\tilde{Q}_{1}(r,\eta)$ for $0\leq r\leq 2$ and $0.7\leq\eta\leq 1$.
-Moreover, $\lim_{r\to 0}Q_{1}(r,\eta;-)>0$ and $\lim_{r\to 0}\tilde{Q}_{1}(r,\eta)=0$ hold.
+7(a), we note that $P_{\mbox{\scriptsize click},1}(r,\eta;-)$ is larger than $\tilde{P}_{\mbox{\scriptsize click},1}(r,\eta)$ for $0\leq r\leq 2$ and $0.7\leq\eta\leq 1$.
+Moreover, $\lim_{r\to 0}P_{\mbox{\scriptsize click},1}(r,\eta;-)>0$ and $\lim_{r\to 0}\tilde{P}_{\mbox{\scriptsize click},1}(r,\eta)=0$ hold.
 Thus, for most realistic cases, the superposition of oppositely squeezed states is more beneficial to the heralded single-photon emission than the pure two-mode squeezed state.
 Looking at Fig.
-7(b), we note that $E(r,\eta;-)<\tilde{E}(r,\eta)$ for $0\leq r\leq 2$ and $0.7\leq\eta\leq 1$.
+7(b), we note that $P_{\mbox{\scriptsize click,c}}(r,\eta;-)<\tilde{P}_{\mbox{\scriptsize click,c}}(r,\eta)$ for $0\leq r\leq 2$ and $0.7\leq\eta\leq 1$.
 However, for larger $r$ and smaller $\eta$ the difference is small.
@@
 We observe that $g^{(2)}(0)$ is always less for $|r;-\rangle$ in the shown region.
-In particular for $|r;-\rangle$, we notice that $g^{(2)}(0)$ approaches zero as $r\to 0$ while for the two-mode squeezed state it is always greater than zero for $r\geq 0$.
+In particular for $|r;-\rangle$, we notice that $g^{(2)}(0)$ approaches zero as $r\to 0$, while for the two-mode squeezed state it is always greater than zero for $r\geq 0$.
 Thus, for $\eta=0.9$ and $0\leq r\leq 0.504$, the single-photon emitter realized by $|r;-\rangle$ has a better quality than the two-mode squeezed state.
@@
 The state $|r;-\rangle$ is a superposition of photon-number states, $|2\rangle$, $|6\rangle$, $|10\rangle$, ….
-Thus, dividing the state $|r;-\rangle$ on two ports of a 50-50 beam splitter, the probability that the imperfect photon detector detects the state $|2\rangle_{a}$ is supressed.
+Thus, dividing the state $|r;-\rangle$ on two ports of a 50-50 beam splitter, the probability that the imperfect photon detector detects the state $|2\rangle_{a}$ is suppressed.
 Thus, $g^{(2)}(0)$ hardly suffers from an imperfection caused by the term $\eta(1-\eta)|2\rangle_{a}{}_{a}\langle 2|$ included in the observable $\hat{M}$ given by Eq.
@@
 This fact implies that the heralded single-photon emitter realized by $|r;-\rangle$ is more robust than that realized by $\hat{S}_{ab}(ir)|0\rangle_{a}|0\rangle_{b}$ under the condition that we must use the imperfect single-photon detector.
+In Fig.
+9(a), we note that $g^{(2)}(0)$ of the pure two-mode squeezed state exhibits partial stability around $g^{(2)}(0)=0.22$ for $0\leq r\leq 1.0$.
+In contrast, our proposed method can achieve a high quality of single photons for the small $r$ regime ($0\leq r\leq 0.504$).
+This regime of the squeezing parameter $r$ is most relevant in the current experiment.
+To compare our scheme with other single-photon sources, the method of the pure two-mode squeezed state in particular, we also need to count the probability of single-photon generation as shown in Fig.
+3.
+When $r$ is smaller than $0.504$, the success probability of obtaining a single photon after the postselection is nearly equal to $1/2$, as shown by $P(1,1;r;-)$ in Fig.
+3(b), hence both the success probability and the quality are high enough.
+However, as Fig.
+3(a) indicates, the yielding of the success case is rather low when $r$ approaches zero.
+The usual trade-off between purity and overall efficiency remains in our scheme [44, 45].
 V Conclusion
-In this paper, we proposed a new method to generate a heralded single-photon source created from a superposition of oppositely squeezed states.
-We outlined the method for producing this state by injecting squeezed and coherent light beams onto a cross-Kerr nonlinear medium and estimated the probability and efficiency of the emission of the heralded single photons.
+In this paper, we proposed a method to generate a heralded single-photon source created from a superposition of oppositely squeezed states.
+We outlined the method for producing this state by injecting squeezed and coherent light beams onto a cross-Kerr nonlinear medium and estimated the probability and conditional probability of the emission of the heralded single photons.
 To evaluate the quality of our proposed method, we used the pure two-mode squeezed state for comparison.
-When the single-photon detector detects the heralding signals inaccurately, our method shows better performances for the second-order intensity correlation function $g^{(2)}(0)$ than the method of the pure two-mode squeezed states.
+When the single-photon detector detects the heralding signals inaccurately, our method shows better performances for the second-order intensity correlation function $g^{(2)}(0)$ than the method of the pure two-mode squeezed states for small $r$ as explained in Sec.
+IV.
 This is because $|r\rangle-|-r\rangle$ is a superposition of photon-number states, $|2\rangle$, $|6\rangle$, $|10\rangle$, … and the probability amplitude of $|n\rangle_{a}|2\rangle_{b}$ for $n=0,1,2,...$ is suppressed when we inject it into the 50-50 beam splitter.
-That is an important advantage of scheme.
+That is an important advantage of our scheme.
 However the generation of the odd superposition of squeezed states is more difficult leading to a natural trade-off.